### v1

@@ IV-A3 Cross-Domain Few-Shot Object Detection (CD-FSOD) Benchmark
TABLE I: Ablation study on the Bangladeshi Traffic Sign Detection Dataset (BDTSD) showing the impact of incorporating various modifications into the proposed pipeline. Here, the columns with $Δ\Updelta$ indicate the increase in the mAP relative to the baseline model, demonstrating the contribution of each component to the overall performance.
@@
IV-B Implementation Details
Our proposed framework, FUSED-Net, leverages a variant of Faster R-CNN that is seamlessly integrated within the Detectron2 framework. Specifically, we employ ResNet-101 in combination with a Feature Pyramid Network (FPN) as the backbone of the architecture, ensuring robust multi-scale feature extraction. For optimization, we utilized the standard Stochastic Gradient Descent (SGD) solver with a momentum of $0.9$, accompanied by a weight decay parameter of $10^{-4}$ to prevent overfitting. Data augmentation was a critical aspect of our implementation, and we applied color jitter with an $0.8$ probability using the transformations provided by the torchvision library [58], enhancing the diversity and resilience of the training data.
In our experimental setup, we benchmarked our framework against several state-of-the-art architectures, including TFA w/cos, A-RPN, DeFRCN, FRCN-ft, and CD-FSOD. During the initial trials, we encountered NAN/INF bounding box errors, which arose due to the constraints imposed by our single GPU environment. These issues were linked to the original configurations of the SOTA architectures, which were not directly compatible with our computational setup. The Detectron2 framework provided a solution to this challenge by recommending adjustments to the base learning rate and batch size according to the number of available GPU devices. To ensure that all comparisons were fair and consistent, we standardized the experimental conditions across all architectures, setting the batch size to $2$ and the base learning rate to $0.0025$. Aside from these adjustments, we adhered closely to the standard Faster R-CNN configuration as provided by the Detectron2 framework to implement FUSED-Net. To maintain impartiality in the performance comparison of FUSED-Netwith other SOTA architectures, we have used 11% of the total data available in BDTSD [18] as the query set. We have utilized COCO-style average precision (AP). This measurement is taken over several thresholds, from 0.5 to 0.95, with a step size of 0.05.
For transparency and to facilitate future research, all code related to our implementation, including the modifications necessary for each architecture, will be made publicly available on GitHub upon acceptance of this manuscript.
IV-C Ablation Study
@@
To assess the contribution of the individual components of FUSED-Net, we conducted an ablation study. The study was performed on BDTSD across 1-shot, 3-shot, 5-shot, and 10-shot scenarios. Our model integrates four key modifications: unfrozen parameters, pseudo-support set, embedding normalization, and domain adaptation. The results, presented in Table I, demonstrate the incremental improvements each modification contributes to the overall system performance, measured by mAP.
IV-C1 Unfrozen Parameters
@@
The decision to unfreeze all layers of our architecture during fine-tuning stage played a pivotal role in improving cross-domain detection accuracy. Unlike conventional practices where certain layers are frozen to retain pre-trained knowledge, our approach allows every layer—from the Backbone to the Region Proposal Network, Feature Pyramid Network, and fully connected layers—to adjust its weights. This flexibility is helpful in scenarios where the target domain (e.g., Bangladeshi Traffic Signs) is significantly different from the base domain (e.g., general object detection datasets). As a result, the model can learn new, domain-specific features that are essential for detecting traffic signs under diverse conditions. The ablation study reveals that unfreezing the parameters alone provides a moderate boost in performance, yielding an $11.2$ mAP in the 1-shot scenario, and continues to enhance performance as more shots are introduced.
IV-C2 Pseudo Support Set
@@
IV-C3 Embedding Normalization
Incorporating embedding normalization further refined our model’s ability to handle high intra-class variance. By standardizing the feature representations of individual instances, our approach reduces the impact of outliers and ensures that features are comparable across different instances. This normalization is enabled by a cosine similarity-based classifier, which focuses on the angular relationships between feature vectors, thereby reducing the influence of magnitude differences. The results shown in Table I highlight the significant improvement this modification brings, particularly in 10-shot scenarios, where mAP increased by $4.6\%$ compared to the baseline. This demonstrates the effect of embedding normalization in enhancing the model’s robustness and generalization capabilities.
TABLE II: Comparison of mAP Performance across Different Few-Shot Object Detection Models on the Bangladeshi Traffic Sign Detection Dataset (BDTSD) under 1-Shot, 3-Shot, 5-Shot, and 10-Shot Scenarios. Models are sorted by the average mAP ($\mu$) across all shots, with the best results highlighted in boldface.
Model 1-shot 3-shot 5-shot 10-shot $\mathbf{\mu}$
TFA w/cos [20] 4.22 6.54 8.19 9.16 7.03
A-RPN [27] 8.11 9.18 10.57 12.22 10.02
DeFRCN [28] 12.53 21.93 25.84 32.28 23.15
FRCN-ft [11] 11.36 18.58 31.19 42.12 25.81
CD-FSOD [29] 12.14 19.61 31.02 43.02 26.45
FUSED-Net (Ours) 28.62 42.53 46.06 54.34 42.89
IV-C4 Domain Adaptation
Domain adaptation had the most significant impact on our model’s performance. By training the model on traffic signs from diverse countries, we reduced the distributional variance between the source and target domains. This strategy was crucial in enabling the model to generalize effectively across different environments. The results in Table I underscore the effectiveness of domain adaptation, showing a dramatic increase in mAP across all shot scenarios. Even in the 1-shot scenario, the mAP jumps by $18.2\%$ compared to the baseline. It remains substantial across higher-shot scenarios, demonstrating the effectiveness of this technique in addressing the challenges posed by domain discrepancies.
IV-D Performance comparison with the State-of-the-Art in Few-Shot Traffic Sign Detection
As shown in Table II, the performance of our proposed architecture was benchmarked against several state-of-the-art FSOD models on BDTSD across 1-shot, 3-shot, 5-shot, and 10-shot scenarios.
@@ IV-D Performance comparison with the State-of-the-Art in Few-Shot Traffic Sign Detection
Our proposed architecture, integrating the strengths of the state-of-the-art approaches while addressing their limitation, consistently outperformed all other models across every shot scenario, achieving the highest average mAP $42.89$. Specifically, the inclusion of a cosine-similarity-based classifier, inspired by TFA w/cos, enhances embedding normalization and comparability across instances, significantly improving performance in few-shot scenarios. Additionally, we incorporate domain adaptation strategies to ensure robust generalization to the target domain. However, unlike the complex methodologies of CD-FSOD, we achieve this through a simpler, more efficient process that is particularly well-suited to the consistent nature of traffic sign detection. Our pseudo-support set technique enriches the training process by introducing necessary variability, crucial in few-shot learning, and ensures that the model is well-calibrated to the specific characteristics of BDTSD.
IV-E Qualitative Analysis
For the qualitative analysis, we evaluated our model’s performance by fine-tuning it on the MTSD dataset and testing it with a 5-shot scenario from BDTSD. In Figs. 5 and 6, the green bounding boxes represent detections made by our model, while the blue bounding boxes denote the ground truth annotations. The labels above the blue boxes indicate the actual class labels, whereas the labels above the green boxes indicate the class labels predicted by our model. Note that, only detected boxes with confidence scores of $0.5$ or higher are displayed.
As demonstrated in 5a, our model effectively detects traffic signs that are blurry and located at significant distances from the camera, a challenging scenario that often leads to detection failures in other models. Additionally, 5b shows that our model can accurately detect and classify multiple traffic signs within a single image, even when they are positioned at varying distances and embedded in complex backgrounds. This ability to handle both single and multiple instances under challenging conditions underscores the robustness of our approach in real-world scenarios.
@@ IV-E Qualitative Analysis
TABLE III: Comparison of mAP Performance Across Various Few-Shot Object Detection Models on the Cross-domain Few-shot Object Detection (CD-FSOD) Benchmark under 1-Shot, 5-Shot, and 10-Shot Scenarios across Three Datasets: Arthropod Taxonomy Orders Object Detection Dataset (ArTaxOr), Underwater Object Detection Dataset (UODD), and Dataset for Object Detection in Aerial Images (DIOR). Models are sorted by the average mAP ($\mu$) across all shots and datasets, with the best results highlighted in boldface.
@@
IV-F Performance Analysis on Cross-Domain Few-Shot Object Detection (CD-FSOD) Benchmark
To thoroughly assess the cross-domain performance of our proposed architecture, we conducted extensive experiments on the CD-FSOD benchmark, which evaluates models using base datasets from one domain (MS COCO) and novel datasets from distinct domains (ArTaxOr, UODD, DIOR). The results of these experiments, as shown in Table III, offer insights into the efficacy of our approach compared to the state-of-the-art (SOTA) FSOD models.
As discussed before, the traditional FSOD benchmarks, such as MS COCO and Pascal VOC, typically involve dividing data into base and novel sets drawn from the same domain. However, this does not reflect real-world applications where domain shifts are common. For instance, a model trained on natural images might need to detect objects in aerial or optical remote sensing images—domains with significantly different characteristics. Models fine-tuned within the same domain may perform well under such circumstances but often struggle when applied across domains. To identify whether it can excel in cross-domain settings, our proposed model, while originally intended for few-shot traffic sign detection, was subjected to rigorous testing on the CD-FSOD benchmark to evaluate its versatility. The results across 1-shot, 5-shot, and 10-shot scenarios on the ArTaxOr, UODD, and DIOR datasets consistently demonstrate our model’s superior performance. Notably, our model achieves the highest average mAP across all datasets and shot settings ($14.2$), outperforming the leading SOTA models, including CD-FSOD which introduced the benchmark.
@@ IV-F Performance Analysis on Cross-Domain Few-Shot Object Detection (CD-FSOD) Benchmark
In conclusion, our model FUSED-Net demonstrates outstanding performance in cross-domain FSOD, surpassing current SOTA architectures in terms of average mAP. This achievement, coupled with the model’s simplicity and robustness, positions our approach as a highly effective solution for real-world FSOD tasks, even in scenarios involving significant domain shifts.

### v2

@@ IV-A1 Datasets
IV-A2 Implementation Details
Framework and Architecture
Our proposed framework, FUSED-Net, leverages a variant of Faster R-CNN seamlessly integrated within the Detectron2 framework. Specifically, we employ ResNet-101 in combination with a Feature Pyramid Network (FPN) as the backbone, ensuring robust multi-scale feature extraction. Data augmentation was a critical aspect of our implementation; we applied color jitter with a probability of $0.8$ (as recommended by [52]) using the transformations provided by the torchvision library [60], enhancing the diversity and resilience of the training data.
TABLE I: Comparison of mAP Performance across Different Few-Shot Object Detection Models on the Bangladeshi Traffic Sign Detection Dataset (BDTSD) under 1-Shot, 3-Shot, 5-Shot, and 10-Shot Scenarios. Models are sorted by whether the architecture has RPN or not, and then by the average mAP ($\mu$) across all shots, with the best results highlighted in boldface.
Model 1-shot 3-shot 5-shot 10-shot $\mathbf{\mu}$
One-Stage Detectors
DETR-ft [10] 2.00 6.90 8.10 13.98 7.75
YOLOv8-ft [9] 12.80 21.80 31.20 45.50 27.83
Two-Stage Detectors
TFA w/cos [22] 4.22 6.54 8.19 9.16 7.03
A-RPN [29] 8.11 9.18 10.57 12.22 10.02
DeFRCN [30] 12.53 21.93 25.84 32.28 23.15
FRCN-ft [13] 11.36 18.58 31.19 42.12 25.81
CD-FSOD [31] 12.14 19.61 31.02 43.02 26.45
FUSED-Net (Ours) 26.64 44.22 45.94 54.44 42.81
Optimization Strategy
We utilized the Stochastic Gradient Descent (SGD) solver with a momentum of $0.9$ and a weight decay parameter of $10^{-4}$ to prevent overfitting. During the initial trials, we encountered NAN/INF bounding box errors due to optimization instability caused by the mismatch between the high default learning rate and the low batch size required by our single GPU environment. The small batch size of $2$ introduced significant gradient noise, and the high default learning rate ($0.02$) exacerbated this issue, producing overly aggressive updates that led to unstable bounding box predictions and NAN/INF errors during loss computation.
To address this, we followed Detectron2 recommendations and adjusted the base learning rate and batch size to suit our computational setup. Specifically, we reduced the batch size to $2$ to avoid memory errors and recalibrated the learning rate to $0.0025$ using the formula provided in the Detectron2 documentation [^2] https://github.com/facebookresearch/detectron2/issues/1128. This adjustment stabilized the training process and ensured consistent convergence across all architectures, including our proposed one.
Hyperparameter Choices
In our experiments, we utilized a scaling factor, $\gamma$, of $20$ for the cosine-similarity-based classifier. This decision was guided by the findings of [22], which demonstrated that a scaling factor of $20$ achieves an effective balance between reducing intra-class variance and maintaining high detection accuracy for novel classes. This choice minimizes the drop in detection accuracy for base classes, a critical consideration in few-shot scenarios with limited training examples.
Benchmarks and Metrics
We benchmarked our framework against several state-of-the-art architectures, including TFA w/cos, A-RPN, DeFRCN, FRCN-ft, and CD-FSOD. To ensure fairness and consistency, we utilized 11% of the total data available in BDTSD [20] as the query set, as suggested by [22]. We employed COCO-style average precision (AP) as our evaluation metric, measured over thresholds ranging from $0.5$ to $0.95$ with a step size of $0.05$.
Reproducibility
To maintain transparency and facilitate future research, all code related to our implementation will be made publicly available on GitHub upon acceptance of this manuscript.
IV-B Model Evaluation and Analysis
IV-B1 Performance comparison with the State-of-the-Art in Few-Shot Traffic Sign Detection
As shown in Table I, the performance of our proposed architecture was benchmarked against several state-of-the-art FSOD models on BDTSD dataset across 1-shot, 3-shot, 5-shot, and 10-shot scenarios.
We evaluated DETR-ft and YOLOv8-ft, two advanced architectures representing single-stage detectors on the BDTSD dataset. However, these models underperformed in few-shot scenarios. DETR-ft achieved an average mAP of only $7.74$, struggling particularly in the lower-shot conditions due to its heavy reliance on large datasets and prolonged training to optimize its complex attention mechanisms. Similarly, YOLOv8-ft attained an average mAP of $27.83$, surpassing CD-FSOD but failing to compete with FUSED-Net. While YOLOv8-ft demonstrated reasonable performance, its lack of a dedicated region proposal network, crucial for localizing objects in data-scarce environments, limits its effectiveness in FSOD tasks. Hence, we moved on to two-stage detectors.
TABLE II: Comparison of mAP Performance across Different Categories of the Bangladeshi Traffic Sign Detection Dataset (BDTSD) under 1-shot, 3-shot, 5-shot, and 10-shot Scenarios. The categories are sorted alphabetically.
Category 1-shot 3-shot 5-shot 10-shot $\mathbf{\mu}$
Cross Roads 16.94 49.84 35.32 63.11 41.30
Narrow Bridge 32.56 38.15 53.64 77.73 50.52
No Overtaking 15.55 58.39 56.99 56.83 46.94
No Use Of Horn 28.28 35.09 55.22 41.14 39.93
Pedestrain Crossing 38.49 41.08 52.90 61.93 48.60
Road Hump 32.17 44.95 45.25 57.69 45.01
School 37.22 45.81 53.29 59.89 49.05
Sharp Bend To The Left 28.23 40.05 36.66 42.96 36.97
Sharp Bend To The Right 19.44 19.33 14.73 33.20 21.67
Side Road Left 25.44 38.21 29.50 49.78 35.73
Side Road Right 36.33 46.20 53.29 48.86 46.17
Speed Limit 40 km/h 33.23 54.35 55.49 52.50 48.89
Speed Limit 60 km/h 54.13 33.29 67.55 73.00 56.99
Speed Limit 80 km/h 27.04 54.20 56.95 64.63 50.70
U Turn 4.34 39.00 32.22 31.88 26.86
Our proposed architecture, integrating the strengths of the state-of-the-art approaches while addressing their limitation, consistently outperformed all other models across every shot scenario, achieving the highest average mAP $42.81$. Specifically, the inclusion of a cosine-similarity-based classifier, inspired by TFA w/cos, enhances embedding normalization and comparability across instances, significantly improving performance in few-shot scenarios. Additionally, we incorporate domain adaptation strategies to ensure robust generalization to the target domain. However, unlike the complex methodologies of CD-FSOD, we achieve this through a simpler, more efficient process that is particularly well-suited to the consistent nature of traffic sign detection. Our pseudo-support set technique enriches the training process by introducing necessary variability, crucial in few-shot learning, and ensures that the model is well-calibrated to the specific characteristics of BDTSD.
IV-B2 Per-Category Performance Analysis
To provide a more granular evaluation of FUSED-Net, we analyze its performance across individual traffic sign categories in the BDTSD dataset. Table II reports the mAP values for each of the 15 categories under 1-shot, 3-shot, 5-shot, and 10-shot scenarios.
The analysis reveals notable variablity in performance across different categories. For instance, the model achieves high average mAP for “Speed Limit 60 km/h” (56.99) and “Narrow Bridge” (50.52), indicating the effective detection in these scenarios. However, the performance for certain categories like “U Turn” (26.68) and “Sharp Bend to the Right” (21.67) is relatively lower, suggesting that these categories present greater challenges for few-shot detection. The underperformance in the “U Turn” and “Sharp Bend to the Right” categories can be attributed to a combination of factors. These include geometric complexity and ambiguity, where subtle distinctions in directional arrows are challenging to capture with limited training data, leading to potential misclassifications. Additionally, the minimalist visual features of these signs, often lacking text or intricate patterns, make them harder to distinguish, especially under conditions like occlusion or poor lighting. Few-shot models face inherent challenges in fine-grained detection tasks, as limited samples restrict their ability to generalize nuanced differences, further compounded by category distribution imbalances. Lastly, these signs often depend on contextual cues from the road environment, which are difficult to incorporate effectively with minimal training data.
The analysis underscores that while FUSED-Net demonstrates robust performance in most categories, it also highlights the impact of category-specific characteristics on detection accuracy. This evaluation provides further insight into the strengths and potential areas of improvement for FUSED-Net, particularly in challenging traffic sign categories.
IV-B3 Statistical Significance of the Performance
Considering the substantial performance gap of FUSED-Net over other SOTA architectures, we conducted multiple independent trials for each of the 1-shot, 3-shot, 5-shot, and 10-shot scenarios to ensure the reproducibility and validity of our findings. After performing the trials, we calculated the mean and $95\%$ confidence intervals (CI95) for the mAP values across different few-shot scenarios. The values are shown in Fig. 6. The results demonstrate a consistent performance advantage of FUSED-Net compared to SOTA methods. The relatively narrow confidence intervals affirm the stability and reliability of the proposed architecture under various few-shot scenarios.
Figure 6: Average mAP across different shot scenarios with 95% confidence intervals. Each bar represents the mean mAP for a given shot scenario, with error bars indicating the range of variability.
To provide a more comprehensive evaluation, we also analyzed the precision, recall, and F1 scores achieved by FUSED-Net as summarized in Table III. The precision and recall values indicate that the model maintains an effective balance between detecting relevant instances (recall) and minimizing false positives (precision) across all shot scenarios. Notably, recall exhibits slightly higher values than precision, suggesting that FUSED-Net prioritizes capturing more potential traffic sign instances, even if it occasionally predicts additional false positives. The F1-score, representing the harmonic mean of precision and recall, improves consistently as the number of shots increases, demonstrating the model’s ability to generalize better with additional data.
TABLE III: Average precision, recall, and F1 scores of FUSED-Net under 1-shot, 3-shot, 5-shot, and 10-shot scenarios.
Modality 1-shot 3-shot 5-shot 10-shot
Precision 0.29 0.43 0.47 0.54
Recall 0.46 0.57 0.58 0.63
F1-Score 0.35 0.49 0.52 0.58
To further validate the significance of these findings, we performed a statistical significance analysis using paired t-tests [61] to compare FUSED-Net’s performance with the next-best-performing SOTA method (CD-FSOD). Across all scenarios, a t-statistic of 5.73 and a p-value of 0.0106 confirm that the observed performance differences are statistically significant.
The narrow confidence intervals, coupled with the precision, recall, and F1-score analysis and the statistical significance of our findings, demonstrate the high reliability of FUSED-Net’s performance. These results instill confidence in the practical applicability of our approach, particularly in real-world scenarios where data availability is often limited.
IV-C Model Insights and Behavior
IV-C1 Ablation Study
@@
TABLE IV: Ablation study on the Bangladeshi Traffic Sign Detection Dataset (BDTSD) showing the impact of incorporating various modifications into the proposed pipeline. Here, the columns with $Δ\Updelta$ indicate the increase in the mAP relative to the baseline model, demonstrating the contribution of each component to the overall performance.
@@
To assess the contribution of the individual components of FUSED-Net, we conducted an ablation study. The study was performed on BDTSD across 1-shot, 3-shot, 5-shot, and 10-shot scenarios. Our model integrates four key modifications: unfrozen parameters, pseudo-support set, embedding normalization, and domain adaptation. The results, presented in Table IV, demonstrate the incremental improvements each modification contributes to the overall system performance, measured by mAP.
Unfrozen Parameters
@@
The decision to unfreeze all layers of our architecture during fine-tuning stage played a pivotal role in improving cross-domain detection accuracy. Unlike conventional practices where certain layers are frozen to retain pre-trained knowledge, our approach allows every layer—from the Backbone to the Region Proposal Network, Feature Pyramid Network, and fully connected layers—to adjust its weights. This flexibility is helpful in scenarios where the target domain (e.g., Bangladeshi Traffic Signs) is significantly different from the base domain (e.g., general object detection datasets). As a result, the model can learn new, domain-specific features that are essential for detecting traffic signs under diverse conditions.
While the ablation study reveals that unfreezing parameters alone yields a moderate improvement of $0.8\%$ mAP in the 1-shot scenario, this result aligns with findings from prior research [31], which demonstrated that unfrozen parameter yields varying degrees of improvement across different FSOD architectures (as illustrated in Fig. 1). For example, DeFRCN, a SOTA FSOD architecture known for its generalizability, achieved only a 1.2% improvement with unfrozen parameters. Similarly, our proposed architecture, designed for high generalizability, exhibits smaller yet meaningful gains, particularly under the highly challenging 1-shot setting.
Further, the impact of unfrozen parameters becomes more apparent in the 5-shot scenario, where a 2.1% improvement in mAP is observed. This trend suggests that while the benefits of unfreezing are modest in extremely low-data regimes, they become increasingly effective as more training data becomes available. Overall, these results demonstrate that unfreezing parameters contributes meaningfully to FUSED-Net’s adaptability and performance, particularly as the difficulty of the task or the availability of data increases.
Pseudo Support Set
@@
To validate our choice of Color Jittering, we conducted additional experiments comparing the effects of various augmentation strategies such as random cropping and horizontal flipping — both individually and in combination with Color Jitter. The results demonstrated that employing Color Jitter alone consistently produced the highest mAP scores across all shot scenarios. These findings reinforce that the simplicity and effectiveness of Color Jitter make it the best choice for augmenting the PSS.
Embedding Normalization
Incorporating embedding normalization further refined our model’s ability to handle high intra-class variance. By standardizing the feature representations of individual instances, our approach reduces the impact of outliers and ensures that features are comparable across different instances. This normalization is enabled by a cosine similarity-based classifier, which focuses on the angular relationships between feature vectors, thereby reducing the influence of magnitude differences. The results shown in Table IV highlight the significant improvement this modification brings, particularly in 10-shot scenarios, where mAP increased by $4.6\%$ compared to the baseline. This demonstrates the effect of embedding normalization in enhancing the model’s robustness and generalization capabilities.
@@
Domain Adaptation
Domain adaptation had the most significant impact on our model’s performance. By training the model on traffic signs from diverse countries, we reduced the distributional variance between the source and target domains. This strategy was crucial in enabling the model to generalize effectively across different environments. The results in Table IV underscore the effectiveness of domain adaptation, showing a dramatic increase in mAP across all shot scenarios. Even in the 1-shot scenario, the mAP jumps by $16.2\%$ compared to the baseline. It remains substantial across higher-shot scenarios, demonstrating the effectiveness of this technique in addressing the challenges posed by domain discrepancies.
IV-C2 Qualitative Analysis
For the qualitative analysis, we evaluated our model’s performance by fine-tuning it on the MTSD dataset and testing it with a 5-shot scenario from BDTSD. In Figs. 7 and 8, the green bounding boxes represent detections made by our model, while the blue bounding boxes denote the ground truth annotations. The labels above the blue boxes indicate the actual class labels, whereas the labels above the green boxes indicate the class labels predicted by our model. Note that, only detected boxes with confidence scores of $0.5$ or higher are displayed.
As demonstrated in 7a, our model effectively detects traffic signs that are blurry and located at significant distances from the camera, a challenging scenario that often leads to detection failures in other models. Additionally, 7b shows that our model can accurately detect and classify multiple traffic signs within a single image, even when they are positioned at varying distances and embedded in complex backgrounds. This ability to handle both single and multiple instances under challenging conditions underscores the robustness of our approach in real-world scenarios.
@@ IV-C2 Qualitative Analysis
IV-D Performance Analysis on Cross-Domain Few-Shot Object Detection (CD-FSOD) Benchmark
To thoroughly assess the cross-domain performance of our proposed architecture, we conducted extensive experiments on the CD-FSOD benchmark, which evaluates models using base datasets from one domain (MS COCO) and novel datasets from distinct domains (ArTaxOr, UODD, DIOR). The results of these experiments, as shown in Table V, offer insights into the efficacy of our approach compared to the state-of-the-art (SOTA) FSOD models.
TABLE V: Comparison of mAP Performance Across Various Few-Shot Object Detection Models on the Cross-domain Few-shot Object Detection (CD-FSOD) Benchmark under 1-Shot, 5-Shot, and 10-Shot Scenarios across Three Datasets: Arthropod Taxonomy Orders Object Detection Dataset (ArTaxOr), Underwater Object Detection Dataset (UODD), and Dataset for Object Detection in Aerial Images (DIOR). Models are sorted by the average mAP ($\mu$) across all shots and datasets, with the best results highlighted in boldface.
@@ IV-D Performance Analysis on Cross-Domain Few-Shot Object Detection (CD-FSOD) Benchmark
As discussed before, the traditional FSOD benchmarks, such as MS COCO and Pascal VOC, typically involve dividing data into base and novel sets drawn from the same domain. However, this does not reflect real-world applications where domain shifts are common. For instance, a model trained on natural images might need to detect objects in aerial or optical remote sensing images—domains with significantly different characteristics. Models fine-tuned within the same domain may perform well under such circumstances but often struggle when applied across domains. To identify whether it can excel in cross-domain settings, our proposed model, while originally intended for few-shot traffic sign detection, was subjected to rigorous testing on the CD-FSOD benchmark to evaluate its versatility.
While our model achieves the highest average mAP across all datasets and shot settings ($14.2$), outperforming the leading SOTA models, including CD-FSOD, it exhibits inconsistencies in the UODD dataset at the 10-shot scenario ($12.7\%$ vs CD-FSOD’s $14.5\%$). This discrepancy likely stems from differences between traffic sign images and underwater object images. Underwater scenes often feature clear, uniform backgrounds (e.g., blue ocean water) and slightly blurry objects, whereas traffic scenes contain highly cluttered and noisy backgrounds.
The Pseudo Support Set technique used in FUSED-Net is highly effective for challenging scenarios like traffic signs with complex surroundings, enabling it to detect such objects robustly. However, in simpler environments like underwater scenes, this method may inadvertently cause overfitting, which limits its adaptability. Despite these domain-specific challenges, our model outperforms SOTA models in overall cross-domain performance and consistently excels in 5-shot scenarios, emphasizing its utility in diverse applications.
@@
In conclusion, while FUSED-Net demonstrates slight domain-specific limitations in simpler datasets like UODD, the overall results firmly establish it as a versatile and robust architecture for cross-domain FSOD. This achievement, coupled with its simplicity and robustness, positions our approach as a highly effective solution for real-world FSOD, even in scenarios involving significant domain shifts.